# Towards Swift Serverless LLM Cold Starts with ParaServe

### v1

@@ Abstract
With the surge in number of large language models (LLMs), the industry turns to serverless computing for LLM inference serving. However, serverless LLM serving suffers from significant cold start latency and service level objective (SLO) violations due to the substantial model size, which leads to prolonged model fetching time from remote storage. We present ParaServe, a serverless LLM serving system that minimizes cold start latency through the novel use of pipeline parallelism. Our insight is that by distributing model parameters across multiple GPU servers, we can utilize their aggregated network bandwidth to concurrently fetch different parts of the model. ParaServe adopts a two-level hierarchical design. At the cluster level, ParaServe determines the optimal degree of parallelism based on user SLOs and carefully places GPU workers across servers to reduce network interference. At the worker level, ParaServe overlaps model fetching, loading, and runtime initialization to further accelerate cold starts. Additionally, ParaServe introduces pipeline consolidation, which merges parallel groups back to individual workers to maintain optimal performance for warm requests. Our comprehensive evaluations under diverse settings demonstrate that ParaServe reduces the cold start latency by up to 4.7 and improves SLO attainment by up to 1.74 compared to baselines.
@@ 1 Introduction
Large language models (LLMs) have become an essential part of daily work and life, powering a wide range of applications such as chatbots [1, 2, 3], programming assistants [4, 5, 6], and AI-enhanced search engines [7, 8]. With the widespread adoption of LLMs, the industry has released numerous LLMs with varying structures, sizes, or training corpora for diverse scenarios [9, 10, 1, 2, 3, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Businesses of different scales are also seeking to leverage custom LLMs tailored to their unique needs [20, 21, 22, 23], rather than relying on general-purpose LLM APIs from AI service providers.
@@
Serverless computing [24, 25] has become a compelling choice for users to deploy their LLM serving pipelines [26, 27, 28, 29]. In serverless LLM serving, users only need to store the model weights in cloud storage and upload an image that runs LLM services. The cloud platform automatically provisions resources to instantiate the LLM, and dynamically scales the number of workers based on the actual request load. This approach offers significant cost savings through pay-per-use billing, where users are only charged for the resource consumption during active request processing.
@@
However, realizing the serverless paradigm for LLM serving is challenging, primarily due to the long latency of cold starts, i.e., provisioning new workers when there’s not enough to meet the current load. Such cold starts are common in serverless workloads due to their bursty traffic patterns, as shown in public traces [30]. As cold starts take place on the critical path of request handling, they directly impact service level objectives (SLOs). The impact is especially pronounced for LLMs that range in size from gigabytes to terabytes. For instance, in our production environment, a cold-start instance produces the first token after up to 40 seconds, whereas subsequent generation takes only 30ms per token. This contrast stresses the need for faster cold starts.
In serverless serving, the LLMs are retrieved on demand from remote model registry, which we refer to as model fetching. Serverless functions have limited network bandwidth in nature. As LLMs are large, model fetching costs significant time that contributes most to the total cold start latency. Prior works utilize caching to alleviate the model fetching time [27, 31]. The idea is to cache LLMs in memory or SSDs to accelerate setting up a GPU worker. While caching is effective for the most frequently accessed models, the problem remains unsolved for the long-tail customers, who derive the most benefit from the serverless paradigm [24].
We present ParaServe, a serverless LLM serving system designed to minimize cold start latency. The insight of ParaServe is simple yet effective—whereas a single server has limited bandwidth, we can distribute the workers across different servers, with each fetching a part of the model in parallel. It is viable because LLMs follow a layered structure; the layers are accessed sequentially, obviating the need to load the entire model at once. ParaServe can start the inference once the first few layers are ready on the first server, and forward the intermediate results to subsequent servers that host the rest of the layers. This design, known as pipeline parallelism, has been well studied in distributed model training and inference [32, 33, 34, 35, 36, 37, 38]. ParaServe applies pipeline parallelism in a novel way to reduce the blocking effect of model fetching.
We take a two-level hierarchical approach to realize the parallel model fetching in our production environment. At the cluster level, we design an algorithm to compute the appropriate parallelism size (i.e., the number of servers to distribute the model) based on the model size, cluster resource usage, and user SLOs. We then use a network-contention-aware worker placement policy to choose a set of servers to host the pipeline. At the worker level, we overlap remote-to-host model fetching, host-to-GPU model loading, and the initialization of container and GPU runtime to further reduce the cold start latency.
@@
Although pipeline parallelism can effectively reduce the cold start latency for the initial inference, it incurs inevitable overheads in subsequent rounds due to the transmissions of intermediate results across workers. To address this problem, we propose pipeline consolidation. Specifically, we allow workers to fetch and load the remaining parts of the model in the background, eventually transitioning back to local inference with fully-loaded model. Depending on the load, each participant worker can independently decide whether to release its resources or load the full model after cold starts. Therefore, with a single cold start, ParaServe can scale up as many workers as the pipeline size, allowing for faster elasticity.
@@
We design and implement ParaServe, a serverless LLM serving system that leverages pipeline parallelism to reduce cold start latency.
We adopt a hierarchical approach that combines cluster-level and worker-level optimizations to satisfy user SLOs with pipeline parallelism.
We propose pipeline consolidation for faster scaling of LLM workers without extra cold starts.
Experiments on real-world datasets show that ParaServe reduces the cold start latency by up to 4.7 and improves SLO attainment by up to 1.74 compared to baselines.
@@ 2 Background and Motivation
In this section, we introduce large language models (LLMs) and serverless LLM serving, and identify opportunities to reduce serverless cold start latency.
@@ 2.1 LLM inference
An LLM takes as input a sequence of tokens, called prompt, and generates an output sequence in an autoregressive fashion, meaning that it generates one token at a time based on both the prompt and its previous outputs. The core component of an LLM is the self-attention layer, which needs to compute the key, value, and query vectors of tokens in the input sequence. Since the key and value vectors of previous tokens remain unchanged during iterations, LLM serving systems usually cache these vectors to avoid redundant computation, known as KV cache. The stage that generates the first token is called prefill, during which the key and value vectors of all tokens in the prompt are calculated and stored in GPU memory. Subsequent tokens are generated in the decoding phase, which reuses the key-value cache and generates one token at a time.
Pipeline parallelism is a model parallelism strategy that is commonly used in LLM training to enable the scaling of model sizes beyond the capacity of a single GPU [32, 33, 34, 35, 36, 37]. In this method, different layers of the model are distributed across multiple workers and inter-worker communication is needed to transfer intermediate results. To minimize pipeline bubbles, a mini-batch of training examples is usually split into micro-batches that are executed in pipeline [32, 34]. Pipeline parallelism introduces little communication overhead between workers. For example, running an input with 256 tokens using OPT-175B generates intermediate messages of 6MB, which can be transmitted within single-digit milliseconds over a 10Gbps network.
2.2 Serverless LLM Serving
In serverless LLM serving, users upload (1) LLM models to a model registry, and (2) a function image containing the serving framework and runtime libraries to a container image registry. The serving framework is responsible for fetching the model from the remote registry, loading the model into GPU memory, and running the model to generate tokens upon user requests. The serverless platform automatically adjust the number of function workers to match the current load. One of the most attractive offerings of serverless LLM serving is its pay-per-use billing, where users are charged only for the running period of each worker and not for the idle time. Therefore, serverless is cost-efficient for long-tail customers whose workloads are sporadic and highly unpredictable.
@@ 2.2 Serverless LLM Serving
In LLM serving, users often specify service level objectives (SLOs) that represent performance expectations [39, 40, 41, 38]. These SLOs are typically made up of two metrics: time to first token (TTFT) and time per output token (TPOT). TTFT represents the time elapsed from user sending the request to the generation of the first token, while TPOT is the average time taken to generate each subsequent token. Different workloads prioritize these metrics differently. For example, real-time chatbots [1, 2, 3] emphasize low TTFT, whereas tasks like article writing prefer lower TPOT [39].
@@
Despite the potential benefits of serverless LLM serving, existing systems suffer from significant cold start latency. A cold start occurs when an incoming request finds no available worker hosting the target model, requiring the system to initialize a new worker. Figure 1 provides a breakdown of cold start latency in our production serverless LLM serving platform, using Llama2-7B model on the NVIDIA A10 GPU. In general, a cold start involves the following stages:
@@
CUDA Context Initialization. The Python runtime set up the CUDA context to prepare for GPU tasks.
@@
Model Loading. The fetched model is loaded into GPU memory.
Inference. The serving framework runs the model to generate the first token of the input sequence.
As show in Figure 1, it takes up to 40 seconds to produce the first token for a Llama2-7B model, whereas each subsequent token is generated in 30 milliseconds on average. Model fetching is the most time-consuming among all stages since the network bandwidth of inference servers are limited and the model weights are large. Furthermore, network contention can occur when multiple cold start containers are placed on the same server. They compete for the network bandwidth, leading to extremely long fetching latency. Container creation also costs a considerable amount of time, as LLM workloads require large function images.
2.3 Opportunies
@@
This paper employs a two-level hierarchical approach to achieve maximum overlapping in the cold start critical path for serverless LLM serving. At the cluster level, we leverage pipeline parallelism for coarse-grained overlapping of model fetching across workers, addressing the most time-consuming phase of cold starts. Upon cold start, we create a pipeline parallelism group across GPU servers, with each worker hosting a part of the model. At the worker level, we perform fine-grained overlapping by carefully reorganizing the intra-worker workflow in Figure 1. First, we offload model fetching from user container to a system-level service, allowing the fetching process to begin before container creation. Second, we notice that model and library loading stages are bound by different resources: model loading is GPU-bound, while library loading is CPU-bound. Thus, we run the two stages in parallel. We swap the order of library loading and CUDA context initialization, and trigger model loading once CUDA context has been prepared. Finally, we pipeline the fetching and loading at tensor granularity to hide the loading overhead. After all initialization stages have finished, the worker runs the model and generates the first token, completing the cold start. The optimized cold-start workflow is shown in Figure 2.
3 ParaServe Overview
ParaServe is a serverless LLM serving system designed to minimize the cold start latency. Unlike traditional serverless model serving systems that load the entire model in a single worker, ParaServe leverages parallel model fetching to significantly reduce model fetching time.
@@ 3 ParaServe Overview
As shown in Figure 3, ParaServe consists of two optimization levels. At the cluster level, the central controller employs a parallelism size selection algorithm (§4.1) and a network-contention-aware worker placement strategy (§4.2) to reduce cold start latency through pipeline parallelism. The parallelism size selection algorithm determines the appropriate pipeline parallelism size for each cold-start model based on user SLOs and cluster resource availability. For a given parallelism size, ParaServe partitions the model across multiple workers and assigns these workers to different GPU servers using the network-contention-aware worker placement strategy. This strategy identifies potential network contentions on GPU servers and ensures that such contention does not cause SLO violations.
@@
At the worker level, ParaServe applies fine-grained overlapping across cold start stages to reduce worker initialization time (§5). First, a node-level model prefetcher is used to proactively fetches model weights for cold-start workers, overlapping model fetching with container creation and runtime initialization time. Second, ParaServe prioritizes the CUDA context initialization and introduces a parameter manager that loads parameters to GPU in parallel with library loading. Additionally, model fetching and loading are pipelined to hide the model loading overhead. These overlapping strategies reduce the overhead of cold start stages other than model fetching, thereby highlighting the value of pipeline parallelism.
@@
Finally, as pipeline parallelism can degrade inference performance, ParaServe performs pipeline consolidation (§6) to transform pipeline workers into standalone workers that host all parameters. ParaServe allows some of cold-start workers to continue loading the remaining model parts while serving requests. Eventually, these workers evolve into individual workers that host the entire model. Each worker independently decides whether to load additional parameters or terminate after completing the cold start process. Ongoing requests, initially processed in the pipeline parallelism group, are migrated to the standalone workers.

### v2

# HydraServe: Minimizing Cold Start Latency for Serverless LLM Serving in Public Clouds

@@ Abstract
With the proliferation of large language model (LLM) variants, developers are turning to serverless computing for cost-efficient LLM deployment. However, public cloud providers often struggle to provide performance guarantees for serverless LLM serving due to significant cold start latency caused by substantial model sizes and complex runtime dependencies. To address this problem, we present HydraServe, a serverless LLM serving system designed to minimize cold start latency in public clouds. HydraServe proactively distributes models across servers to quickly fetch them, and overlaps cold-start stages within workers to reduce startup latency. Additionally, HydraServe strategically places workers across GPUs to avoid network contention among cold-start instances. To minimize resource consumption during cold starts, HydraServe further introduces pipeline consolidation that can merge groups of workers into individual serving endpoints. Our comprehensive evaluations under diverse settings demonstrate that HydraServe reduces the cold start latency by 1.7– 4.7 and improves service level objective attainment by 1.43–1.74 compared to baselines.
@@ 1 Introduction
Large language models (LLMs) have become an essential part of daily work and life, powering a wide range of applications such as chatbots [1, 2, 3], programming assistants [4, 5, 6], and AI-enhanced search engines [7, 8]. While the industry is releasing numerous LLMs with varying structures, sizes, or training corpora for diverse scenarios [9, 10, 1, 2, 3, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], businesses of different scales are also seeking to develop custom LLMs tailored to their unique needs [21, 22, 23, 24]. As a result, a global ecosystem comprising over one million models has emerged, as documented by Hugging Face [25].
@@
To efficiently deploy these models, serverless LLM serving has emerged as a compelling choice [25, 26, 27, 28, 29, 30, 31, 32, 33]. In serverless inference, the cloud platform automatically provisions GPUs for models based on the actual request load, and model owners are only charged for the resource consumption during request processing. This approach offers significant cost savings for long-tail models with infrequent usage.
@@
However, for cloud providers, delivering performance guarantees for serverless LLM users is challenging, primarily due to the long latency of cold starts, i.e., provisioning new workers to meet the current load. Production traces reveal that bursty traffic patterns are prevalent in both LLM serving workloads [34] and serverless workloads [35], making cold starts inevitable in serverless LLM serving [30, 32, 31, 33]. Cold starts can directly impact service level objectives (SLOs) for LLM inference. For instance, in our production environment, a cold-start instance produces the first token after more than 40 seconds, whereas subsequent generation takes only 30ms per token. This contrast stresses the need for faster cold starts.
Addressing this problem is non-trivial, as LLM cold starts introduce two unique challenges compared to traditional serverless workloads. First, LLM applications require massive external state, with model weights fetched on demand from remote registries during cold starts. Since serverless instances typically operate with constrained network bandwidth, model fetching becomes the primary bottleneck in cold start latency. Second, the LLM runtime is highly complex, involving multiple specialized libraries including CUDA runtime [36], AI frameworks [37, 38], and inference engines [39, 40, 41]. The complex dependencies of these libraries lead to substantial container creation and library loading overhead [42]. In public clouds, applications run in isolated containers with varying library versions, making it impractical to prepare runtimes in advance for every user.
Prior works focus on reducing the model fetching latency and propose two approaches: (1) caching LLMs in memory or SSDs [30, 33, 43] and (2) retrieving model weights from active workers holding the same model [31, 32, 33]. While effective for frequently accessed models, these optimizations fall short for long-tail customers, who benefit the most from the serverless paradigm [44]. Additionally, retrieving data from peer workers requires high-bandwidth networking, which is not cost-efficient for serverless inference clusters.
To this end, we present HydraServe, a serverless LLM serving system designed to minimize cold start latency in public clouds. HydraServe incorporates two key insights to mitigate the overhead of model fetching and runtime preparation. First, while a single server has limited bandwidth for model fetching, distributing workers across multiple servers enables efficient bandwidth aggregation. Second, since model fetching and runtime preparation exhibit weak execution dependencies, these stages can be parallelized rather than executed sequentially, thereby overlapping their latencies.
Specifically, to distribute models across servers, we take advantage of the layered structure of LLMs. HydraServe partitions LLM layers across servers and exchanges intermediate results during inference. This design, known as pipeline parallelism, has been well studied in distributed model training and inference [45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Pipeline parallelism is also used to quickly start inference when scaling up multiple workers [32]. HydraServe further adopts pipeline parallelism to reduce the blocking effect of model fetching—we proactively create a group of workers in parallel even if the model only needs one.
@@
Although pipeline parallelism can effectively reduce the cold start latency for the initial inference, it may impact performance in subsequent rounds due to GPU sharing. To address this, we propose pipeline consolidation. Specifically, we allow workers to fetch and load the layers previously assigned to other workers in the background, eventually transitioning back to local inference with a fully-loaded model. Depending on the load, each worker can independently decide whether to release its resources or load the full model after cold starts. Therefore, with a single cold start, HydraServe can choose to only create a single worker in the end or scale up as many workers as the pipeline size, allowing for high elasticity.
HydraServe takes a three-level hierarchical approach to realize the above optimizations. At the cluster level, we design an algorithm to allocate resources for cold-start models based on the model size, cluster resource usage, and user SLOs. A network-contention-aware worker placement policy is adopted to avoid potential network contention. At the worker level, we overlap remote-to-host model fetching, host-to-GPU model loading, and the initialization of container and GPU runtime to further reduce the cold start latency. Finally, at the inference level, we consolidate workers that are previously created in pipeline groups and migrate their runtime states to ensure peak performance with all weights loaded.
@@
We articulate the main bottleneck of LLM cold starts in public clouds and identify opportunities to mitigate the overhead.
We propose a hierarchical approach that combines cluster-level, worker-level, and inference-level optimizations to reduce cold start latency for serverless LLM serving in public clouds.
We implement HydraServe in both testbed and production environment. Experiments on real-world datasets show that HydraServe reduces the cold start latency by 1.7– 4.7 and improves SLO attainment by 1.43–1.74 compared to baselines. The brownfield evaluation further confirms these gains, yielding an average latency reduction of 2.6.
@@ 2 Background and Motivation
This section provides an overview of large language models (LLMs) and serverless LLM serving. We highlight the critical issue of limited network bandwidth for LLM serving in public clouds, which can substantially increase cold start latency. We then identify opportunities for mitigating this performance bottleneck in a cost-effective manner.
@@ 2.1 LLM Inference
An LLM takes as input a sequence of tokens, called prompt, and generates an output sequence in an autoregressive manner, producing one token at a time based on both the prompt and previously generated outputs. The core component of an LLM is the self-attention layer, which needs to compute the key, value, and query vectors of tokens in the input sequence. Since the key and value vectors of previous tokens remain unchanged during iterations, LLM serving systems usually cache these vectors to avoid redundant computation, known as KV cache. The stage that generates the first token is called prefill, during which the key and value vectors of all tokens in the prompt are calculated and stored in GPU memory. Subsequent tokens are generated in the decoding phase, which reuses the key-value cache and generates one token at a time.
@@
In LLM serving, users often specify service level objectives (SLOs) that represent performance expectations [55, 56, 57, 51]. These SLOs typically focus on two metrics: time to first token (TTFT) and time per output token (TPOT). TTFT measures the latency from request submission to the generation of the first token, while TPOT represents the average time taken to generate each subsequent token. The relative importance of these metrics varies across different applications. For example, real-time chatbots emphasize low TTFT, whereas article writing prefers lower TPOT [55].
Pipeline parallelism is a model parallelism strategy that is commonly used in LLM training to scale up involved GPUs [45, 46, 47, 48, 49, 50, 58]. This approach distributes a model’s layers across multiple workers, with intermediate results transmitted sequentially between workers during computation.
2.2 Serverless LLM Serving in Public Clouds
| Instance | Mem.(GB) | Band.(Gbps) | #GPU | Cost($/h) | Cost/GPU($/h) |
| g6e.xlarge | 32 | up to 20 | 1 | 1.861 | 1.861 |
| g6e.2xlarge | 64 | up to 20 | 1 | 2.24208 | 2.24208 |
| g6e.4xlarge | 128 | 20 | 1 | 3.00424 | 3.00424 |
| g6e.8xlarge | 256 | 25 | 1 | 4.52856 | 4.52856 |
| g6e.16xlarge | 512 | 35 | 1 | 7.57719 | 7.57719 |
| g6e.12xlarge | 384 | 100 | 4 | 10.49264 | 2.62316 |
| g6e.24xlarge | 768 | 200 | 4 | 15.06559 | 3.76640 |
| g6e.48xlarge | 1536 | 400 | 8 | 30.13118 | 3.76640 |
Recently, serverless inference has gained widespread adoption among cloud providers, including Aliyun [27], AWS [26], and Azure [59]. In serverless LLM serving, users upload (1) model weights and (2) an image containing the serving framework and runtime dependencies to a cloud storage. This serving framework is responsible for fetching model weights from remote storage, loading them into GPU memory, and running inference upon user requests. The serverless platform automatically scales model workers to accommodate fluctuating loads, including the ability to scale down to zero workers when idle. One of the most attractive offerings of serverless LLM serving is its pay-per-use billing, where users are charged only for the running periods of each worker. In production environments, long-tail models typically exhibit sporadic and unpredictable workload patterns [60], making them ideal candidates for serverless serving [30].
Network bandwidth constraints. Contrary to conventional wisdom, serverless LLM inference may operate with substantially lower network capacity than commonly anticipated. This constraint arises partly from multiple instances sharing a single server, but the primary driver is the cost-efficiency requirement of serverless computing. LLM inference performance mainly depends on the GPU capabilities. Since serverless customers prioritize cost savings, cloud providers aim to minimize cost per GPU when configuring their infrastructure. This cost-optimization strategy favors servers with reduced CPU, memory, and network resources.
Table 1 illustrates this economic principle by comparing configurations and costs of AWS EC2 instance types equipped with NVIDIA L40S GPUs [61]. As Table 1 reveals, resources other than GPUs constitute a significant portion of server costs. Using single-GPU instances as an example, compared to the instance type with the lowest cost per GPU (g6e.xlarge), adding extra resources can increase costs by 20% to 300%. Consequently, servers with constrained non-GPU resources deliver substantial cost savings, making them the preferred choice for serverless providers. However, these economical instances impose network bandwidth limitations that significantly increase cold start latency. While a network-only upgrade is less expensive, the additional network bandwidth would be severely underutilized, as it is only used during cold starts for model fetching. This paper aims to address this trade-off by minimizing cold start latency while preserving the cost benefits of resource-constrained servers.
@@ 2.2 Serverless LLM Serving in Public Clouds
Breakdown of LLM cold starts. Despite the potential benefits of serverless LLM serving, existing systems suffer from significant cold start latency. A cold start occurs when an incoming request finds no available worker hosting the target model, necessitating the creation of a new worker. Figure 1 provides a detailed breakdown of cold start latency in our public serverless inference platform [27], using vLLM [40] to run a Llama2-7B model on NVIDIA A10 GPUs. In general, a cold start involves the following stages:
@@
CUDA Context Initialization. The runtime initializes the CUDA context to prepare for GPU tasks.
@@
Model Loading. The fetched model is loaded into GPU memory and model states including CUDA graph and KV cache are initialized.
Inference. Run the model and generate the first token.
As shown in Figure 1, a cold start requires more than 40 seconds to produce the first token. Model fetching dominates this latency due to limited network bandwidth and contention among colocated containers. Additionally, container creation incurs substantial overhead due to the large image sizes of LLM workloads (8.31 GB in our setup). For model loading, prior work has largely eliminated CUDA graph and KV cache initialization overhead through state materialization [63]. The remaining bottleneck lies in transferring model weights to the GPU, which takes approximately 2 seconds in our setup.
2.3 Opportunities
@@
This paper achieves maximum overlapping to reduce cold start latencies. At the cluster level, we leverage pipeline parallelism for coarse-grained overlapping of model fetching across workers. Upon a cold start, we create a pipeline parallelism group across GPU servers, with each worker only hosting a part of the model. This approach can significantly reduce the single-worker startup latency. At the worker level, we perform fine-grained overlapping by carefully reorganizing the intra-worker workflow in Figure 1. We offload model fetching to a system-level service, allowing fetching to begin before container creation. Furthermore, we observe that model loading (GPU-bound) and library loading (CPU-bound) use different resources, allowing parallel execution. We reorder operations to initialize CUDA context first, and then simultaneously start to load model and library. Finally, we pipeline the fetching and loading at tensor granularity to hide the loading overhead. Figure 2 shows this optimized workflow.
3 HydraServe Overview
@@
HydraServe is a serverless LLM serving system designed to minimize the cold start latency. As shown in Figure 3, HydraServe adopts a three-level hierarchical architecture. At the cluster level, the central controller employs a resource allocation algorithm (§4.1) and a network-contention-aware worker placement strategy (§4.2) for parallel model fetching. The algorithm allocates resources for each cold-start model based on user SLOs, while the placement strategy identifies potential network contentions on GPU servers and ensures that such contention does not cause SLO violations. HydraServe also strives to distribute pipeline workers across GPUs to mitigate performance degradation caused by GPU sharing.
@@
At the worker level, HydraServe applies fine-grained overlapping across cold-start stages to accelerate worker startup (§5). First, a node-level model prefetcher is used to proactively fetch model weights for cold-start workers, overlapping model fetching with container creation and runtime initialization times. Second, HydraServe prioritizes the CUDA context initialization and introduces a parameter manager that loads parameters to the GPU in parallel with library loading. Additionally, model fetching and loading are pipelined to hide the model loading overhead. These overlapping strategies reduce the overhead of cold start stages other than model fetching, thereby highlighting the value of pipeline parallelism.
@@
Finally, at the inference level, HydraServe consolidates pipeline workers into standalone workers that host all parameters (§6). Specifically, HydraServe allows a cold-start worker to continue loading the remaining model parts while serving requests, eventually evolving into an individual worker that hosts the entire model. Each worker can independently decide whether to evolve or terminate after completing the cold start process. Ongoing requests in the pipeline parallelism group are then migrated to the standalone workers.
HydraServe ensures that cold starts incur no higher first-token latencies or request completion times than standard cold-start processes. This guarantee in request completion time is achieved through HydraServe’s request migration mechanism, which ensures peak inference speed after all model weights have been prepared.